1 + cos 0 + sino 1 + cos 0 - sin sec 0 + tan 0
i need to prove the identity
Answer #1

we have the left hand side as

\frac{1+\cos \theta+\sin \theta}{1+\cos \theta-\sin \theta}

we multiply numerator and denominator by 1+\sin \theta

=\frac{1+\cos \theta+\sin \theta}{1+\cos \theta-\sin \theta}*\frac{1+sin \theta}{1+\sin \theta}

=\frac{(1+\cos\theta+\sin \theta)(1+\sin \theta)}{1+\cos \theta-\sin \theta+\sin \theta+\sin \theta\cos \theta-\sin^2\theta}

now using identity

1-\sin^2\theta=\cos^2\theta

=\frac{(1+\cos\theta+\sin \theta)(1+\sin \theta)}{\cos^2\theta+\cos \theta+\sin \theta\cos \theta}

=\frac{(1+\cos\theta+\sin \theta)(1+\sin \theta)}{\cos\theta(1+\cos \theta+\sin \theta)}

=\frac{1+\sin \theta}{\cos \theta}

=\frac{1}{\cos \theta}+\frac{\sin \theta}{\cos \theta}

=\boldsymbol{\sec \theta+\tan \theta}

hence proved ,